the initial rate of hydrolysis of methyj acetate(1M) by a weak acid (HA, 1M) is 1/100th of that of a strong acid ( HX, 1M) at 25 degree celcius. The Ka of HA is

Grade:11

2 Answers

Sunil Kumar FP
10 years ago
given rate(HA)/(HX)=1/100
k(ester)(h+) of[ HA]/k(ester)(h+) of [HX]=.01
[H+] of ha / [H+] of hx =.01
h+ of ha=c×£^2
therefore c×£^2=.01(c=1)(alsp conc of h+ from hx=1)
£=.1

now kA=c×£^2/1-£
kA=10^(-4)

thanks and regards
sunil kumar
askIItian faculty

Sunil Kumar FP
10 years ago
ka=10(-4)

thanks and regards
sunil kr
askIItian faculty